# A Bayesian hierarchical mixture cure modelling framework to utilize multiple survival datasets for long-term survivorship estimates: A case study from previously untreated metastatic melanoma

### v1

@@ 5 Discussion
We have presented a Bayesian hierarchical mixture cure model to obtain complete survival curves which exhibit plateau behaviour. These curves can then be used in HTA to inform functions of patient lifetime. We have demonstrated the novel method using the CheckMate 067 trial data set for end-points OS and PFS and at different artificial data-cuts. The hierarchical model was compared with the separate model analogue and has been shown to be superior, as well as both principled and applicable to situations where trials include multiple treatment arms and end-points.
@@
In many cases, our approach is preferable to a frequentist or non-hierarchical model. A benefit to adopting a Bayesian paradigm over a frequentist approach is the flexibility to modelling complex structure, in our particular case multi-level or nested data. In contrast, a frequentist model would need to impose assumptions about the structure of a covariance matrix. Another advantage of adopting a hierarchical structure in a Bayesian approach is robustness to smaller sample sizes due to the provision of prior information and borrowing of strength between data for different end-points, leading to more stable, accurate extrapolation estimates. Principled prior distribution specification about the cure fractions was shown to be especially useful early on in the trial when observable data about long-term survivors was weak. This results in the stabilisation of the cure fraction estimates, which again results in improved extrapolation. In general, extrapolations of survival outcomes with a shared cure fraction provide a more conservative and practical approach to modelling the end-points separately. It may be that there is excessive borrowing of information between cure fractions when the end-points are distinctly different. In our example, this was not necessary, since there are only two end-points but for a larger number this may be appropriate. Partial-exchangeability can be used instead in this case [26].
@@
In comparison to some previous frequentist approaches, our framework is able to include contextual information about the parameters of the model via the prior distributions. Incorporating sensible information, both in the form of external data sources or elicited from experts can stabilise the inference on the cure fraction and restrict/constrain its values via the prior distributions. For the analysis presented in this paper, we exploited different priors. Minimally informative priors allow the data to dominate the posterior even if the dataset is relatively small. In practice, there may be additional information from clinical experience or previous data that can better inform the prior selection. The priors can be specified on the natural scale to aid elicitation and interpretation. For example, we may have clinical knowledge that a drug treatment could generate a cure fraction greater than, say, 30%. If a (cancer) patient enters the study at 60 years old then there is a high degree of certainty that they will not live for, say, over 40 years. In the case of sparse data e.g. when there is a large amount of missing data perhaps due to right censoring, then ’regularising’ the inference based on available prior information could make a significant and important difference. In particular, the method of blending curves is a recent, straightforward option for how to constrain survival [27]. How to elicit this prior information in practice may not be simple and a formal protocol should be adopted [28].
@@
In our analysis, we simulated censoring due to data-cuts by simply censoring all patients in the complete data set at the same time point. In practice, a more realistic censoring time may have been at the date of last visit for each patient, but this would have made no differences for the purposes of this analysis. Primarily, the data-cut points were selected to demonstrate the theoretical justification of our method and not to strictly follow the study data-cut protocol.
Future work should test the proposed hierarchical MCM on a variety of data sets and at data-cuts reflecting their actual trial protocols. The modelling framework is generalisable from two end-points in our example to any number in principle. For example, previous related frequentist analyses considered clusters of hospitals and recurrent events per individual [10, 9] or interspecies extrapolation in dose-response experiments [29]. Further work could investigate the benefit to including additional end-point data even if it does not clearly demonstrate plateauing behaviour. The background survival curve used in the model representing the statistically ‘cured’ survival in general population can be refined to be appropriate to specific cohort characteristics. For instance, long-term survival of clinically identified complete responders (CR) could be used to bridge the gap between statistical cure and clinical cure/functional cure. The shorter-term CR individual-level survival data can be combined with longer-term external data, such as from WHO life tables as in the current analysis (see [30] for potential approaches).
Another way to think about the cut-point data is to think about what is the minimum follow-up time should we choose to halt the trial and end the data collection which would maximise some expected overall utility. That is, put another way, what is the optimal stopping time $\tau^{*}$ for the total utility $V$?
@@
The stopping utility should incorporate a measure of the survival extrapolation robustness and may include benefits, such as smaller costs, freed-up time from the trial for patients and those running the study, quicker time to market for a drug, and avoidance of side-effects etc. The utility would also depend on estimating ‘correct’ parameters values from the available data ie, that they are within some acceptable threshold of the true values. This could be applied directly on the parameters of interest or an order statistics for which drug is deemed preferable.
This may be framed in terms of value of information (VoI) [31]. For example, in the case given above, what is the additional value of obtaining further data after 12 and 30 months. In practice, deriving a set of heuristics may be better than formalising a general rule. A discrete set of cut-points may be used, similar to our artificial selection of rounded times 12 and 30 months, rather than considering the whole real line.
@@
There are several implications for HTA by using the hierarchical mixture cure model presented here. In a cost-effectiveness analysis, unstable estimates over time would give different estimates of benefit and thus cost-effectiveness statistics and potentially differing optimal decisions. Lee (2019)[32] performed a cost-effectiveness analysis of using nivolumab with ipilimumab vs ipilimumab using the CheckMate 067 trial data, and a partitioned survival model and Markov state-transition model. Lifetime costs and benefits were estimated. Results were obtained using 18-month (when OS data were unavailable) and 36-month (OS available) CheckMate 067 data-cuts. They showed that the model using OS data generated more than 1 additional quality-adjusted life-year (QALY) across both treatment arms compared to the model without OS data. Using our approach, we have shown that reliable estimates of complete survival curves for OS can be obtained even for early data-cuts. These can be used to estimate lifetime horizon benefits in instances where perhaps previously estimates were not available. In our example, the OS cure fraction is accurate and stable at 12 months for the hierarchical mixture cure model with exponential distributions for the OS and PFS uncured survival models. Additional benefits are potential early end of trial (cheaper, improved health, quicker drugs to market for producers and patients benefit), and smaller trials.

### v2

@@ 6 Discussion
We have presented a Bayesian hierarchical mixture cure model to obtain complete survival curves that exhibit plateau behaviour. These curves can then be used in HTA to inform functions of patient lifetime. We also conducted a simulation study demonstrating situations where the hierarchical or separate models are more appropriate. Finally, we demonstrated the novel method using the CheckMate 067 trial data set for end-points OS and PFS and at different artificial data-cuts; incorporating additional structure for this particular situation. The hierarchical model was compared with the separate model analogue and has been shown to be superior, as well as both principled and applicable to situations where trials include multiple treatment arms and end-points.
@@
Our approach may be preferable to a frequentist or non-hierarchical model both practically, as shown in the simulation study, and conceptually, given the theoretical framework underpinning it. A benefit to adopting a Bayesian paradigm over a frequentist approach is the flexibility in modelling complex structure, in our particular case multi-level or nested data. In contrast, a frequentist model would need to impose assumptions about the structure of a covariance matrix. Another advantage of adopting a hierarchical structure in a Bayesian approach is robustness to smaller sample sizes due to the provision of prior information and borrowing of strength between data for different end-points, leading to more stable, accurate extrapolation estimates. Principled prior distribution specification about the cure fractions was shown to be especially useful early on in the trial when observable data about long-term survivors was weak. This results in the stabilisation of the cure fraction estimates, which again results in improved extrapolation. In general, extrapolations of survival outcomes with a shared cure fraction provide a more conservative and practical approach to modelling the end-points separately. It may be that there is excessive borrowing of information between cure fractions when the end-points are distinctly different. In our example, this was not necessary, since there are only two end-points but for a larger number this may be appropriate. Partial-exchangeability can be used instead in this case 28.
@@
In comparison to some previous frequentist approaches, our framework is able to include contextual information about the parameters of the model via the prior distributions. Incorporating sensible information, both in the form of external data sources or elicited from experts can stabilise the inference on the cure fraction and restrict/constrain its values via the prior distributions. For the analysis presented in this paper, we exploited different priors. Minimally informative priors allow the data to dominate the posterior, even when the dataset is relatively small. In practice, there may be additional information from clinical experience or previous data that can better inform the prior selection. The priors can be specified on the natural scale to aid elicitation and interpretation. For example, we may have clinical knowledge that a drug treatment could generate a cure fraction greater than, say, 30%. If a (cancer) patient enters the study at 60 years old then there is a high degree of certainty that they will not live for, say, over 40 years. In the case of sparse data e.g. when there is a large amount of missing data perhaps due to right censoring, then ‘regularising’ the inference based on available prior information could make a significant and important difference. In particular, the method of blending curves is a recent, straightforward option for how to constrain survival 29. How to elicit this prior information in practice may not be simple and a formal protocol should be adopted 30.
@@
In our analysis, we simulated censoring due to data-cuts by simply censoring all patients in the complete data set at the same time point. In practice, a more realistic censoring time might have been the date of last visit for each patient; however, this would not have affected the results of this analysis. Primarily, the data-cut points were selected to demonstrate the theoretical justification of our method and not to strictly follow the study data-cut protocol.
Future work should test the proposed hierarchical MCM on a variety of data sets and at data-cuts reflecting their actual trial protocols. The modelling framework is generalisable from two end-points in our example to any number in principle. For example, previous related frequentist analyses considered clusters of hospitals and recurrent events per individual 10, 9 or interspecies extrapolation in dose-response experiments 31. Further work could investigate the benefit to including additional end-point data, even if it does not clearly demonstrate plateauing behaviour. The background survival curve used in the model representing the statistically ‘cured’ survival in general population can be refined to be appropriate to specific cohort characteristics. For instance, long-term survival of clinically identified complete responders (CR) could be used to bridge the gap between statistical cure and clinical cure/functional cure. The shorter-term CR individual-level survival data can be combined with longer-term external data, such as from WHO life tables as in the current analysis (see 32 for potential approaches). As a way to model the constraint that $S_{OS}\geq S_{PFS}$ we could consider an excess hazard-type formulation such as $h_{OS}=h_{B}+h_{OSe}$ and $h_{PFS}=h_{B}+h_{OSe}+h_{PFSe}$, where $h_{OS}$ is the hazard for all cause mortality, $h_{PFS}$ is the hazards for either progression or all cause mortality, $h_{B}$ is the background hazard, $h_{OSe}$ is the excess hazard for overall survival and $h_{PFSe}$ is the additional excess survival due to progression alone.
Another way to think about the cut-point data is to reflect on what is the minimum follow-up time should we choose to halt the trial and end the data collection which would maximise some expected overall utility. That is, put another way, what is the optimal stopping time $\tau^{*}$ for the total utility $V$?
@@
The stopping utility should incorporate a measure of the survival extrapolation robustness and may include benefits, such as smaller costs, freed-up time from the trial for patients and those running the study, quicker time to market for a drug, and avoidance of side-effects. The utility would also depend on estimating ‘correct’ parameters values from the available data ie, that they are within some acceptable threshold of the true values. This could be applied directly on the parameters of interest or an order statistics for which drug is deemed preferable.
This may be framed in terms of value of information (VoI) 33. For example, in the case given above, what is the additional value of obtaining further data after 12 and 30 months. In practice, deriving a set of heuristics may be better than formalising a general rule. A discrete set of cut-points may be used, similar to our artificial selection of rounded times 12 and 30 months, rather than considering the whole real line.
@@
There are several implications for HTA by using the hierarchical mixture cure model presented here. In a cost-effectiveness analysis, unstable estimates over time would give different estimates of benefit and thus cost-effectiveness statistics and potentially differing optimal decisions. Lee (2019)34 performed a cost-effectiveness analysis of using nivolumab with ipilimumab vs ipilimumab using the CheckMate 067 trial data, and a partitioned survival model and Markov state-transition model. Lifetime costs and benefits were estimated. Results were obtained using 18-month (when OS data were unavailable) and 36-month (OS available) CheckMate 067 data-cuts. They showed that the model using OS data generated more than 1 additional quality-adjusted life-year (QALY) across both treatment arms compared to the model without OS data. Using our approach, we have shown that reliable estimates of complete survival curves for OS can be obtained even for early data-cuts. These can be used to estimate lifetime horizon benefits in cases where such estimates were previously unavailable. In our example, the OS cure fraction is accurate and stable at 12 months for the hierarchical mixture cure model with exponential distributions for the OS and PFS uncured survival models. Additional benefits are potential early end of trial (cheaper, improved health, quicker drugs to market for producers and patients benefit), and smaller trials.